GlueX TOF Meeting, January 24, 2017

From GlueXWiki
Jump to: navigation, search

GlueX Time-of-Flight Meeting
Tuesday, January 24, 2017
11:00 am EST
JLab: CEBAF Center, Room F326/327


  1. Announcements
  2. Review of minutes from the previous meeting
  3. Run Readiness
    1. FADC Replacement
    2. Online Monitoring
    3. Hot Check-Out
  4. HV scan
  5. Lucite shield items?
  6. NIM Paper (Paul)
  7. Action Item Recap



Talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2016 on the JLab CUE. This directory is accessible from the web at .



  • FSU: Sasha Ostrovidov
  • JLab: Thomas Britton, Brad Cannon, Mark Ito (chair), Simon Taylor, Beni Zihlmann
  • NU: Sean Dobbs


  • The Hall closes on Friday.
  • Beam due in Hall on Monday during the day.

Review of minutes from the previous meeting

We looked over the minutes from January 10. No significant comments were made.

Run Readiness

  • FADC Replacement. Sascha Somov confirms that the broken FADC that Beni identified is indeed bad and Sascha will replace it. After that Beni will ask Sergey Furletov to do a pedestal setting. And after that Beni will look for channels that have missing bits on their DACs and either (a) set the pedestals by hand or (b) replace the units if spares are available.
  • Online Monitoring. We need a routine such that problems like the bad FADC are caught early. Brad has discussed the issue with David Lawrence; there will be an on-going display of the critical monitoring histograms in the counting room. There should also be a set of standard "good" plots with which to compare. Also it would be nice if there was some sort of histogram-viewing activity on the checklist for shift takers. Thomas volunteered to look into these issues.
  • Hot Check-Out. Beni, Brad, Thomas will get together and check things out.

HV scan

Brad's HV setting jobs are done as of this morning. There were some issues getting the latest version of the code from Beni, but those have been sorted. The HV values agree with Beni's within a few percent. These are for the double-ended counters; single-ended HV values are still to come.

Sean asked about resolution and efficiency with the lower voltages has been studied by the group. He showed some plots from the monitoring plots that show poorer resolution now (at 1200 V for all counters) than in the spring, 140 ps vs. 100 ps for pi-, obtained after Beni's most recent calibrations. Beni will try to confirm these results.

Beni repeated a test with a new tube using the laser in the lab. At 1300 V, he can go to 10 MHz without degradation of the pulse amplitude. At 1500 V, this is no longer the case.

Sean showed other plots showing the rate of TOF hit matches with charged tracks. There is clearly a drop off for the half-width counters. Beni things that that is due to degradation in gain for those tubes because of high currents last Spring. A channel-by-channel HV setting should bring those back. Beni also thinks that after this run, we should replace some of these inner tubes.

Sean remarked that until recently, with the uploading of Beni's new constants, the poor TOF timing was holding up calibration, having effects in several other detectors. Getting good TOF constants early should be a priority this time.

Lucite Shield Simulation

Thomas is working on a simulation of rates in the TOF with and without the Lucite shielding that was installed in December. He is almost there. He still needs some details of the geometry, in particular the dimensions of the inner sleeve that fits inside the beam hole. If he sees an effect comparable to that see in data, then we can study how we might optimize the arrangement.